# Multi-role Consensus through LLMs Discussions for Vulnerability Detection

### 21 Mar 2024 v1

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the potential for vulnerability detection, a crucial component of software quality assurance. Despite this progress, most studies have been limited to the perspective of a single role, usually testers, lacking diverse viewpoints from different roles in a typical software development life-cycle, including both developers and testers. To this end, this paper introduces an approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code. Preliminary evaluation of the proposed approach indicates a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
@@ 1 Introduction
As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats. Recently, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [1, 2].
@@
To bridge this research gap, this paper introduces an approach, namely Multi-role Consensus through LLMs Discussions (MuCoLD), for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process. The key idea of MuCoLD is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
2 MuCoLD
Fig.1 shows the workflow of MuCoLD, where a tester and a developer in a code review team attempt to reach a collective consensus on the existence and classification of vulnerabilities in a code segment by discussing with each other according to their own perspectives through a three-stage process.
@@ 2 MuCoLD
Discussion stage: The discussion stage aims to realize an iterative output exchange in attempt to reach a collectively multi-perspective consensus inside the code review team. The tester and the developer, equipped with their unique perspective and initial judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities. During this stage, the tester and the developer repeats a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning. A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains goal-oriented and time-efficient.
@@
Figure 1: Overview of MuCoLD
@@ 3-A Experiment Settings
Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo model. MuCoLD is compared to the result obtained from the single-role LLM approach. Inspired by [1], the following two types of prompts are used to test these two approaches. The basic prompt directly asks the LLM about the presence of a major vulnerability in FC, AE, AU, or PU without additional context. Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [1, 4].
@@
data ability single-role MuCoLD single-role MuCoLD single-role MuCoLD
@@ 3-B Results and Discussions
Table.I summarizes the experiment results, highlighting the improvements through MuCoLD. In average, there is a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score. In terms of computation costs, due to the need for conversation between different roles, it requires a 478% increase in the number of tokens consumed. The experiment results indicates a more obvious improvement in vulnerability detection performance, particularly when the proportion of vulnerable data in the testing dataset is higher. This is likely because the rounds of discussions allows LLMs to explore a broader spectrum of potential vulnerabilities, leading to a higher likelihood of classifying code as vulnerable. Moreover, the imprecise definitions tied to the vulnerability categories, each associated with a range of Common Weakness Enumeration IDs, may also lead more code segments identified as vulnerable.
@@ 4 Conclusion and Future Work
This paper proposes MuCoLD, an approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus. Preliminary evaluation have shown notable improvements in recall rates and F1 scores. Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, and extend its use to multi-class classification and vulnerability patching, broadening insights and proving its versatility.

### 10 Apr 2024 v2

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the potential for vulnerability detection, a crucial component of software quality assurance. Despite this progress, most studies have been limited to the perspective of a single role, usually testers, lacking diverse viewpoints from different roles in a typical software development life-cycle, including both developers and testers. To this end, this paper introduces a multi-role approach to employ LLMs to act as different roles to simulate real-life code review process, engaging in discussions towards a consensus on the existence and classification of vulnerabilities in the code. Preliminary evaluation of the proposed approach indicates a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score.
@@ 1 Introduction
As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2]. Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4, 5].
@@
To bridge this research gap, this paper introduces a novel multi-role approach, namely Multi-role Consensus through LLMs Discussions, for employing LLMs to act as multiple different roles in vulnerability detection to simulate a real-life code review process. The key idea of this approach is to integrate the diverse perspectives of team members with different roles, thereby reaching a well-informed and collective consensus regarding the classification of major vulnerabilities in the given code.
2 Multi-role Approach
@@
Discussion stage: The discussion stage aims to realize an iterative output exchange in attempt to reach a collectively multi-perspective consensus inside the code review team. The tester and the developer, equipped with their unique perspective and judgments, enters a dialectic interaction, aimed at exploring and resolving different opinions on potential vulnerabilities. During this stage, the tester and the developer repeats a ”pose query - deduce response - relay insight” loop, which serves as an incremental prompt, pushing participants to re-evaluate and polish their judgment and reasoning. A maximum depth on discussion rounds is pre-set to prevent the dialogues from devolving into an endless cycle, ensuring that the discussion remains both goal-oriented and time-efficient.
@@
Figure 1: Overview of the multi-role approach
@@ 3-A Experiment Settings
Baseline: All LLMs used in this experiment are instances of the gpt-3.5-turbo model. The proposed multi-role approach is compared to the result obtained from the single-role LLM approach. Inspired by [3], the following two types of prompts are used to test these two approaches. The basic prompt directly asks the LLM about the presence of a major vulnerability in FC, AE, AU, or PU without additional context. Chain-of-thought (CoT), i.e., step-by-step, is also used to analyze the existence of a particular vulnerability.
Metrics: To assess the enhancements in vulnerability detection performance, this study employs three key metrics: the precision rate, the recall rate, and the F1 score [3].
@@
data ability single-role multi-role single-role multi-role single-role multi-role
@@ 3-B Results and Discussions
Table.I summarizes the experiment results, highlighting the improvements through the proposed approach. In average, there is a 4.73% increase in the precision rate, 58.9% increase in the recall rate, and a 28.1% increase in the F1 score. In terms of computation costs, due to the need for conversation between different roles, it requires a 478% increase in the number of tokens consumed. The experiment results indicates a more obvious improvement in vulnerability detection performance, particularly when the proportion of vulnerable data in the testing dataset is higher. This is likely because the rounds of discussions allows LLMs to explore a broader spectrum of potential vulnerabilities, leading to a higher likelihood of classifying code as vulnerable. Moreover, the imprecise definitions tied to the vulnerability categories, each associated with a range of Common Weakness Enumeration IDs, may also lead more code segments identified as vulnerable.
@@ 4 Conclusion and Future Work
This paper proposes a novel approach for improving vulnerability detection with LLMs, where LLMs act as different roles to simulate their interaction in a real-life code review team, discussing the existence and classification of vulnerabilities in code to arrive at a consensus. Preliminary evaluation have shown notable improvements in recall rates and F1 scores. Future work should enhance this approach by integrating more diverse roles, like program managers to reflect the diversity of a real-world code review team, broadening insights and proving its versatility.